OpenStudy (anonymous):

Write an equation of the line the passes through (-5,8) and (-3,-8) a. y = 8x + 32 b. y = 8x - 32 c. y = -8x + 32 d. y = -8x - 32

OpenStudy (dumbcow):

slope formula for 2 points \[m = \frac{y _{2} - y_{1}}{x _{2} - x_{1}} = \frac{-8 - 8}{-3 - (-5)} = \frac{-16}{2} = -8\] point-slope form: \[y - y_{1} = m(x - x_{1})\] \[y - 8 = -8(x-(-5))\] \[y-8 = -8x - 40\] \[y = -8x -32\]